# Hier-SLAM: Scaling-up Semantics in SLAM with a Hierarchically Categorical Gaussian Splatting

### v3

@@ IV-B SLAM Performance
Tracking Accuracy. We present the tracking performance on the Replica [27] and ScanNet [28] datasets in Tab. I and Tab. II, respectively. On the Replica dataset, our proposed method surpasses all current approaches. For the ScanNet dataset, the performance of all methods is lower than on the synthetic dataset due to the noisy, sparse depth sensor input and the limited color image quality caused by motion blur. We evaluate all six sequences, showing that our method performs comparably to state-of-the-art methods [29].
@@
  • First / Second block results are from NVIDIA GeForce RTX 4090 and NVIDIA L40S, respectively.
@@ IV-C Hierarchical semantic understanding
TABLE V: Semantic metric mIoU (%) comparison and Parameter usage (MB) on Replica. Results are highlighted as first, second.
@@
mIoU (%) NIDS-SLAM [23] 82.37 82.45 84.08 76.99 85.94
@@
Figure 4: Visualization of the established semantic 3D map across multiple levels, demonstrating a coarse-to-fine semantic understanding of the complex scene. The bottom of the figure displays localization, mapping, and rendering performance, providing a comprehensive overview of Hier-SLAM’s effectiveness.
@@ IV-D Scaling up capability
To demonstrate the scaling-up capability, we apply our proposed method to the real-world complex dataset, ScanNet [28], which covers up to 550 unique semantic classes. Unlike Replica [27], where the semantic ground truth is synthesized from a global world model and can be considered ideal, the semantic annotations in ScanNet are significantly noisier. Additionally, the dataset features noisy depth sensor inputs and blurred color images, making semantic understanding particularly challenging in this scenes.
Using the flat semantic representation cannot even run successfully due to storage limitations. In contrast, we establish the hierarchical tree with the assistance of LLMs, which guide the compaction of the coding from the original 550 semantic classes to 72 semantic codings, resulting in over 7 times reduction in coding usage. As visualized in Fig. 4, our estimated 3D global semantic map at different levels demonstrates a coarse-to-fine semantic understanding, showcasing our method’s scaling-up capability in handling this complex scene.
@@ V CONCLUSIONS
We present Hier-SLAM, a novel semantic 3D Gaussian Splatting SLAM method with a hierarchical categorical representation, which can generate global 3D semantic map with scaling-up capability and explicit semantic semantic label prediction. Specifically, we propose a novel hierarchical representation to encode semantic message in a compact form, further forming it into 3D Gaussian Splatting, with the assistance of LLMs. Additionally, we present a novel semantic loss for the hierarchical semantic information optimization, including both inter-level and cross-level optimization. Furthermore, we refine the whole SLAM system. Experiments demonstrate that Hier-SLAM surpasses existing dense SLAM methods in both tracking and mapping performance with faster speed, and it can significantly reduces storage and training requirements. The method exhibits exceptional rendering performance, achieving up to 2,000/3,000 FPS with/without semantic information. Most notably, Hi-SLAM showcases its ability to manage complex real-world scenes containing over 500 semantic classes, underlining its robust scalability and effectiveness in large-scale applications.

### v4

@@ IV-B SLAM Performance
Tracking Accuracy. We present the tracking performance on the Replica [27] and ScanNet [28] datasets in Tab. I and Tab. II, respectively. On the Replica dataset, our proposed method surpasses all current approaches. For the ScanNet dataset, the performance of all methods is lower than on the synthetic dataset due to the noisy, sparse depth sensor input and the limited color image quality caused by motion blur. We evaluate all six sequences, showing that our method performs comparably to state-of-the-art methods [9, 29].
@@
  • First & Second block results are from NVIDIA GeForce RTX 4090 and NVIDIA L40S, respectively.
@@ IV-C Hierarchical semantic understanding
TABLE V: Semantic performance mIoU (%) and Parameter usage (MB) on Replica. Results are highlighted as first, second.
@@
mIoU (%) total 102 classes NIDS-SLAM [23] 82.37 82.45 84.08 76.99 85.94
@@
[…]
SemGauss-SLAM [17] 96.34 96.30 95.82 96.51 96.72
SGS-SLAM [16] 92.72 92.95 92.91 92.10 92.90
Hier-SLAM (Ours) 95.58 95.25 95.81 95.73 95.52
@@
  • Ours represents our method with a hierarchical representation, evaluated on a subset of semantic classes, consistent with [16].
  • First & Second block mIou (%) results are evaluated over a total of 102 semantic classes and a subset of classes, respectively.
Figure 4: Visualization of the established semantic 3D map across multiple levels, demonstrating a coarse-to-fine semantic understanding of the complex scene. The bottom of the figure displays localization, mapping, and rendering performance, providing a comprehensive overview.
@@ IV-D Scaling up capability
To demonstrate the scaling-up capability, we apply our proposed method to the real-world complex dataset, ScanNet [28], which covers up to 550 unique semantic classes. Unlike Replica [27], where the semantic ground truth is synthesized from a global world model and can be considered ideal, the semantic annotations in ScanNet are significantly noisier. Additionally, the dataset features noisy depth sensor inputs and blurred color images, making semantic understanding particularly challenging in this scenes. Using the flat semantic representation cannot even run successfully due to storage limitations. In contrast, we establish the hierarchical tree with the assistance of LLMs, which guide the compaction of the coding from the original 550 semantic classes to 72 semantic codings, resulting in over 7 times reduction in coding usage. As visualized in Fig. 4, our estimated 3D global semantic map at different levels demonstrates a coarse-to-fine semantic understanding, showcasing our method’s scaling-up capability in handling this complex scene.
@@ V CONCLUSIONS
We present Hier-SLAM, a novel semantic 3D Gaussian Splatting SLAM method with a hierarchical categorical representation, which can generate explicit global 3D semantic label map with scaling-up capability. Specifically, we introduce a compact hierarchical representation for semantic encoding, integrating it into 3D Gaussian Splatting with LLMs assistance. We further propose a novel semantic loss for optimizing hierarchical semantic information across inter- and cross-levels. Our refined SLAM system achieves superior or on-par performance with existing dense SLAM methods in tracking, mapping, and semantic understanding while running faster and significantly reducing storage. Hier-SLAM delivers exceptional rendering at up to 2,000/3,000 FPS (with/without semantics) and effectively handles complex real-world scenes, demonstrating strong scalability.
APPENDIX
[…]
tableRendering performance PSNR, SSIM, LPIPS on Replica. Best results are highlighted as first, second. Methods Metrics Avg. room0 room1 room2 office0 office1 office2 office3 office4 Visual SLAM NICE-SLAM [29] PSNR $\uparrow$ 24.42 22.12 22.47 24.52 29.07 30.34 19.66 22.23 24.94 SSIM $\uparrow$ 0.809 0.689 0.757 0.814 0.874 0.886 0.797 0.801 0.856 LPIPS $\downarrow$ 0.233 0.330 0.271 0.208 0.229 0.181 0.235 0.209 0.198 Vox-Fusion [31] PSNR $\uparrow$ 24.41 22.39 22.36 23.92 27.79 29.83 20.33 23.47 25.21 SSIM $\uparrow$ 0.801 0.683 0.751 0.798 0.857 0.876 0.794 0.803 0.847 LPIPS $\downarrow$ 0.236 0.303 0.269 0.234 0.241 0.184 0.243 0.213 0.199 Co-SLAM [32] PSNR $\uparrow$ 30.24 27.27 28.45 29.06 34.14 34.87 28.43 28.76 30.91 SSIM $\uparrow$ 0.939 0.910 0.909 0.932 0.961 0.969 0.938 0.941 0.955 LPIPS $\downarrow$ 0.252 0.324 0.294 0.266 0.209 0.196 0.258 0.229 0.236 ESLAM [33] PSNR $\uparrow$ 29.08 25.32 27.77 29.08 33.71 30.20 28.09 28.77 29.71 SSIM $\uparrow$ 0.929 0.875 0.902 0.932 0.960 0.923 0.943 0.948 0.945 LPIPS $\downarrow$ 0.239 0.313 0.298 0.248 0.184 0.228 0.241 0.196 0.204 SplaTAM [9] PSNR $\uparrow$ 34.11 32.86 33.89 35.25 38.26 39.17 31.97 29.70 31.81 SSIM $\uparrow$ 0.968 0.978 0.969 0.979 0.977 0.978 0.969 0.949 0.949 LPIPS $\downarrow$ 0.102 0.072 0.103 0.081 0.092 0.093 0.102 0.121 0.152 Semantic SLAM SNI-SLAM [22] PSNR $\uparrow$ 29.43 25.91 28.17 29.15 31.85 30.34 29.13 28.75 30.97 SSIM $\uparrow$ 0.921 0.884 0.900 0.921 0.935 0.925 0.930 0.932 0.936 LPIPS $\downarrow$ 0.237 0.307 0.292 0.265 0.185 0.211 0.230 0.209 0.198 SGS-SLAM [16] PSNR $\uparrow$ 34.66 32.50 34.25 35.10 38.54 39.20 32.90 32.05 32.75 SSIM $\uparrow$ 0.973 0.976 0.978 0.981 0.984 0.980 0.967 0.966 0.949 LPIPS $\downarrow$ 0.096 0.070 0.094 0.070 0.086 0.087 0.101 0.115 0.148 SemGauss-SLAM [17] PSNR $\uparrow$ 35.03 32.55 33.32 35.15 38.39 39.07 32.11 31.60 35.00 SSIM $\uparrow$ 0.982 0.979 0.970 0.987 0.989 0.972 0.978 0.972 0.978 LPIPS $\downarrow$ 0.062 0.055 0.054 0.045 0.048 0.046 0.069 0.078 0.093 Hier-SLAM (Ours) PSNR $\uparrow$ 35.70 32.83 34.68 36.33 39.75 40.93 33.29 32.48 35.33 SSIM $\uparrow$ 0.980 0.976 0.979 0.987 0.988 0.989 0.975 0.971 0.976 LPIPS $\downarrow$ 0.067 0.060 0.063 0.052 0.050 0.049 0.083 0.081 0.094
I Detailed Results on Rendering Quality
The detailed rendering quality results are presented in Tab. References, demonstrating that our method outperforms state-of-the-art approaches.